OpenStudy (anonymous):

find the curve of intersection of the elliptic cyclinder x^2 +4z^2 = 16 and the plane x+3y-2z = 3

OpenStudy (anonymous):

You have to impose equality: x^2 + 4*z^2 - 16 = x + 3*y - 2*z - 3 So, you have y = 1/3*(x^2 + 4*z^2 - x + 2*z - 13) which is an ellipse